# Inference for Regression with

### 23 Feb 2024 v1

@@ A.2 Proofs for Section 3
$\left|\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]^{-1}\right|\to_{p}0.$ (21)
Now consider the numerator term. We first write
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;(Y_{i}-\hat{\bm{\theta}}_{i}^{T}\bm{\gamma})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\bm{\gamma}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\varepsilon_{i}=:T_{1,n}+T_{2,n}.$
Consider term $T_{1,n}$. By Assumption 2(iii), we have
$\left|T_{1,n}-T_{1,n,a}-T_{1,n,b}\right|\to_{p}0$
where
$T_{1,n,a}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)\bm{\gamma}$
$T_{1,n,b}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i})^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\bm{\gamma}.$
@@
Assumptions 2(i) and 2(ii) together imply that $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}0$, from which it follows that $T_{1,n,a}\to_{p}0$. For term $T_{1,n,b}$ note that
@@
$=\mathbb{E}\left[\mathbb{E}\left[\left.\left(\frac{\mathbf{x}_{i}}{C_{i}}-\mathbf{p}_{i}\right)\left(\frac{\mathbf{x}_{i}}{C_{i}}-\mathbf{p}_{i}\right)^{T}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbb{E}\left[\frac{\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}}{C_{i}}\right]$
@@
$\mathbb{E}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}\right\|^{2}_{F}\right]$ $=\frac{1}{n}\sum_{j=1}^{V}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{X}_{i}\right)_{j,k}^{2}\right]$
$\leq\frac{1}{n}\sum_{j=1}^{V}\sum_{k=1}^{V}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,j}\right)^{2}\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]$
$\leq\frac{1}{n}\sum_{k=1}^{V}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]\leq\frac{1}{n}\mathbb{E}\left[C_{i}^{-1}\right],$
@@
where the second inequality is because $\hat{\mathbf{p}}_{i}$ is in the simplex and the third inequality is by (22). Hence it follows by Chebyshev’s inequality that
@@
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{X}_{i}\right\|_{F}=O_{p}\left(\mathbb{E}\left[C_{i}^{-1}\right]^{1/2}\right),$
@@
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i})^{T}-\sqrt{n}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}\left(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right)^{T}\right]\right|\to_{p}0.$
Finally, by (7), (22), and Assumption 2(ii) we conclude that
$T_{1,n}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right)\bm{\gamma}.$ (23)
Now consider term $T_{2,n}$. Again by Assumption 2(i)-(iii), we have
$\left|T_{2,n}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\varepsilon_{i}\right)\right|\to_{p}0.$
Consider the sum in parentheses. The summands have mean zero (by (4)) and variance
$\mathbb{E}\left[\varepsilon_{i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|\bm{\theta}_{i}\right]\left(\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{p}_{i})-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\right)\right]$
$=\mathbb{E}\left[\varepsilon_{i}^{2}\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right]+\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\varepsilon_{i}^{2}\left(\mathrm{diag}(\mathbf{p}_{i})-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\right].$ (24)
In the above derivation, the first equality uses conditional independence of $Y_{i}$ and $\mathbf{x}_{i}$ given $(C_{i},\bm{\theta}_{i})$, the second uses independence of $Y_{i}$ and $C_{i}$ and the fact that $\mathbf{x}_{i}$ is multinomial, the third uses independence of $C_{i}$ and $(Y_{i},\bm{\theta}_{i})$. Finally, in view of (7), we obtain
$(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\mathbb{E}\left[\varepsilon_{i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\to\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right].$
Note the right-hand side matrix has full rank by Assumption 2(i). Moreover, as $\hat{\mathbf{p}}_{i}$ takes values in the simplex, we have $\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\hat{\mathbf{p}}_{i}\varepsilon_{i}\|\leq\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\||\varepsilon_{i}|$. Hence, for all $t>0$,
$\mathbb{E}\left[\left\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\hat{\mathbf{p}}_{i}\varepsilon_{i}\right\|^{2}\mathbb{I}\left[\left\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\hat{\mathbf{p}}_{i}\varepsilon_{i}\right\|>t\sqrt{n}\right]\right]\to 0$
by Assumption 2(iv). It follows by the Lindeberg–Feller central limit theorem that
$T_{2,n}\to_{d}N\left(\bm{0},\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right).$ (25)
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right].$
To this end, first note that in view of Assumption 2(iii) we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right|\to_{p}0.$
Write
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2})\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}=:T_{3,n}+T_{4,n}.$
First consider term $T_{3,n}$. By (24) we have
$\mathbb{E}[T_{3,n}]\to\mathbf{B}^{T}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}.$
Moreover, for any positive constant $M$ we can write
$T_{3,n}-\mathbb{E}[T_{3,n}]=\frac{1}{n}\sum_{i=1}^{n}\left(\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right)+\frac{1}{n}\sum_{i=1}^{n}\left(\varepsilon_{-,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\varepsilon_{-,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right)\\ =:T_{3,n,a}+T_{3,n,b},$
where $\varepsilon_{+,i}=\varepsilon_{i}\mathbb{I}[|\varepsilon_{i}|\leq M]$ and $\varepsilon_{-,i}=\varepsilon_{i}\mathbb{I}[|\varepsilon_{i}|>M]$. For term $T_{3,n,a}$, note that the summands have mean zero, satisfy
@@
$\left\|\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right\|\leq 2M^{2},$
and
$\left\|\mathbb{E}\left[\left(\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right)^{2}\right]\right\|\leq\left\|\mathbb{E}\left[\left(\varepsilon_{+,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)^{2}\right]\right\|\leq M^{2}\mathbb{E}\left[\varepsilon_{i}^{2}\right],$
because $\|\hat{\mathbf{p}}_{i}\|\leq 1$. Applying Theorem 1.4 of Tropp, (2012), for any $t>0$ we have
$\Pr\left(\|T_{3,n,a}\|>t\right)\leq V\exp\left(\frac{-t^{2}n^{2}/2}{nM^{2}\mathbb{E}\left[\varepsilon_{i}^{2}\right]+2M^{2}tn/3}\right)\to 0$
provided $M^{2}/n\to 0$ as $n\to\infty$. For $T_{3,n,b}$, we have
$\mathbb{E}[\|T_{3,n,b}\|]\leq 2\mathbb{E}\left[\left\|\varepsilon_{-,i}^{2}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right\|\right]\leq 2\mathbb{E}\left[\varepsilon_{i}^{2}\mathbb{I}[|\varepsilon_{i}|>M]\right]\leq 2\frac{\mathbb{E}\left[\varepsilon_{i}^{2+\delta}\mathbb{I}[|\varepsilon_{i}|>M]\right]}{M^{\delta}}\to 0$
as $M\to\infty$ by Assumption 2(iv). In particular, choosing $M=n^{1/4}$ ensures that both $T_{3,n,a}$ and $T_{3,n,b}$ are asymptotically negligible in which case $T_{3,n}\to_{p}\mathbf{B}^{T}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}$ and so

### 24 Mar 2024 v2

@@ A.2 Proofs for Section 3
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right).$
Proof of Lemma 6.
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii), where
$T_{1,n}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
$T_{2,n}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i})^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}.$
@@
Assumption 2(i)-(ii) implies $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. As $\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 1$, it follows that $T_{1,n}\to_{p}\mathbf{0}$. For term $T_{2,n}$, note by Lemma 1 that
@@
$\mathbb{E}\left[\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{X}_{i}\right\|^{2}_{F}\right]=\sum_{j=1}^{V}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{X}_{i}\right)_{j,k}^{2}\right]$ $\leq\sum_{j=1}^{V}\sum_{k=1}^{V}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,j}\right)^{2}\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]$
$\leq\sum_{k=1}^{V}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]\to 0,$
@@
where the second inequality is because $\hat{\mathbf{p}}_{i}$ is in the simplex and the convergence to zero holds in view of (7) and (22). It follows that
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i})^{T}-\sqrt{n}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}\left(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i}\right)^{T}\right]\right\|\to_{p}0.$
We conclude that $T_{2,n}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right)$ by (7), (22), and Assumption 2(i)-(ii)
Lemma 7.
Let Assumption 2(i)-(iv) hold. Then
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}\mathbf{0}.$
Proof of Lemma 7.
@@
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii)-(iv), where
$T_{1,n}$ $=\left(\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
$T_{2,n}$ $=\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i})^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}.$
Assumption 2(i)-(ii) implies that $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. Moreover, $\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|$, which is bounded in probability by Assumption 2(iv). It follows that $T_{1,n}\to_{p}\mathbf{0}$. For $T_{2,n}$, note that $\mathbb{E}[\mathbf{q}_{i}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}]=\mathbf{0}$ by independence of $\mathbf{q}_{i}$ and $\mathbf{x}_{i}$ conditional on $(C_{i},\bm{\theta}_{i})$. Let $\mathbf{X}_{i}=\mathbf{q}_{i}\left(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right)^{T}$ and $D$ the dimension of $\mathbf{q}_{i}$. Then
$\mathbb{E}\left[\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{X}_{i}\right\|^{2}_{F}\right]=\sum_{j=1}^{D}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{X}_{i}\right)_{j,k}^{2}\right]$ $=\sum_{j=1}^{D}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{q}_{i,j}\right)^{2}\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]$
$\leq\sum_{j=1}^{D}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{q}_{i,j}\right)^{4}\right]^{1/2}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{4}\right]^{1/2}$
$\leq\mathrm{constant}\times\sum_{k=1}^{V}\mathbb{E}\left[\left(\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}\right)^{2}\right]^{1/2}\to 0,$
where the first inequality is by Cauchy-Schwarz, the second is by Assumption 2(iv) and the fact that $|\hat{\mathbf{p}}_{i,k}-\mathbf{p}_{i,k}|\leq 1$, and convergence to zero is by (7) and (22). It follows that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{X}_{i}\to_{p}\mathbf{0}$. We conclude by Assumption 2(i)-(ii) that $T_{2,n}\to_{p}\mathbf{0}$. ∎
Lemma 8.
Let Assumption 2(vi) hold. Then
$\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|\to_{p}0.$
Proof of Lemma 8.
Let $\|\,\cdot\,\|_{1}$ be the $\ell^{1}$ norm. As $\hat{\mathbf{p}}_{i}|(C_{i},\bm{\theta}_{i})\sim{C_{i}}^{-1}\mbox{Multinomial}(C_{i},\mathbf{p}_{i})$, for all $t>0$ we have
$\Pr\left(\left.\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1}>t\right|\{(C_{i},\bm{\theta}_{i})\}_{i=1}^{n}\right)\leq\sum_{i=1}^{n}(2^{V}-2)\exp\left\{-\frac{C_{i}t^{2}}{2K}\right\}$
by the union bound and Lemma 1 of Mardia et al., (2019). Then by Assumption 2(vi),
$\Pr\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1}>t\right)\leq n(2^{V}-2)\exp\left\{-\frac{c(\log n)^{1+\epsilon}t^{2}}{2K}\right\},$
where $c,\epsilon>0$. Hence, $\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1}\to_{p}0$. The result now follows because the $\ell^{1}$ norm is weakly greater than the Euclidean norm. ∎
Proof of Theorem 2.
First consider the denominator term. By Lemma 5 and condition (7), we have $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}]$. Hence, by Assumption 2(iv),
$\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}]^{-1}.$ (23)
Now consider the numerator term. We have
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;(Y_{i}-\hat{\bm{\xi}}_{i}^{T}\bm{\psi})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}(\bm{\xi}_{i}-\hat{\bm{\xi}}_{i})^{T}\bm{\psi}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\varepsilon_{i}=:T_{1,n}+T_{2,n}.$
First consider term $T_{1,n}$. By definition,
$T_{1,n}=\left[\begin{array}[]{c}\bm{S}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\right)\bm{S}^{T}\bm{\gamma}\\ \left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\right)\bm{S}^{T}\bm{\gamma}\end{array}\right].$
It follows by Lemmas 6 and 7 that
$T_{1,n}\to_{p}\left[\begin{array}[]{c}-\kappa\,\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$ (24)
Now consider term $T_{2,n}$. By construction,
$T_{2,n}=\left[\begin{array}[]{c}\bm{S}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\varepsilon_{i}\right)\\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}\varepsilon_{i}\end{array}\right].$
We may deduce by arguments similar to those in the proof of Lemma 6 that
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\varepsilon_{i}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\varepsilon_{i}\right)\right\|\to_{p}0.$
under Assumption 2(i)-(iii). Moreover,
$\mathbb{E}\left[\varepsilon_{i}^{2}\left\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right\|^{2}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.\left\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right\|^{2}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|C_{i},\bm{\theta}_{i}\right]\frac{1}{C_{i}}\mathrm{tr}\left\{\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right\}\right]$
$=\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\varepsilon_{i}^{2}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right]\to 0.$
In the above display, the first equality is by independence of $(Y_{i},\mathbf{q}_{i})$ and $\mathbf{x}_{i}$ conditional on $(C_{i},\bm{\theta}_{i})$, the second is by Lemma 1, and the third is by independence of $C_{i}$ and $(Y_{i},\bm{\theta}_{i},\mathbf{q}_{i})$. Therefore,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})\varepsilon_{i}\to_{p}0$
@@
$\left\|T_{2,n}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bm{\xi}_{i}\varepsilon_{i}\right\|\to_{p}0.$
Note that $\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ is finite by Assumption 2(iv)-(v). It follows by the central limit theorem that
$T_{2,n}\to_{d}N\left(\bm{0},\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\right).$ (25)
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right].$
To this end, first write
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\left(\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-{\bm{\xi}}_{i}\!\!\!\;{\bm{\xi}}_{i}^{T}\right)\\ +\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2})\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=:T_{3,n}+T_{4,n}+T_{5,n}.$
@@
Evidently, $T_{3,n}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ by the LLN. Now consider $T_{4,n}$. By construction, we have